# The Representation and Recall of Interwoven Structured Knowledge in LLMs: A Geometric and Layered Analysis

### v1

@@ Appendix C Blessing of dimensionality
When the dimensionality is very high, the most of random vector pairs approach orthogonality. We illustrate this by sampling pairs of vectors in an 8129-dimensional space (corresponding to the activation vector dimension of Meta-Llama-3.1-70B) and computing their cosine similarities. The 99.9% confidence interval (CI) provides an estimate of the expected cosine similarity range at each dimensionality:
@@
Figure C1: Cosine similarity distribution of random vector pairs in an 8129-dimensional space, with a 99.9% confidence interval ($-0.0364$, $0.0364$) shown by the dashed lines.
The specific distribution is shown in Fig.C1, where the confidence interval is extremely narrow (±$\pm 0.036$), indicating that random vector pairs exhibit highly consistent cosine similarities. This suggests that the learned weights of the linear probe across different feature pairs are effectively uncorrelated, exhibiting only random alignment. In Fig.8, the shaded region represents the 99.9% confidence interval for the cosine similarities of high-dimensional random vectors, further supporting this observation.
Appendix D Linear probing detailed results
In Sec.3.1, we applied linear probing to train linear-kernel Support Vector Regression (SVR) models for each layer $l$ and attribute $A_{j}$, using the activation dataset $H_{j}(l)$. The ground truth values $y_{i,j}$ correspond to the attribute values of each element $X_{i}$.
D.1 Why $R^{2}$ cannot reach 1
Even with a perfectly trained model and a sufficiently large dataset, achieving an $R^{2}$ of 1 in linear probing is impossible. The model’s output token ‘1’ indicates that the activation at the last token position in the final layer leads to the highest logit for token 1’s ID after the final linear transformation. However, do token activation in the final layer of numbers exhibit a perfect linear relationship with their real numerical values? Token embeddings are learned representations that capture semantic relationships between tokens, but they are not guaranteed to align linearly with numerical values.
@@ D.1 Why $R^{2}$ cannot reach 1
D.2 Detailed results of the best layer
@@
In the main text Sec.3, we used SVR for linear probing. Figures D1, D2, D3, D4, and D5 present the detailed $R^{2}$ performance of the best layer for each attribute—atomic number, atomic mass, electronegativity, period, and group—across three different models.
@@
Figure D1: Evaluation of SVR performance for Layer best_layer on the atomic mass. The left plot shows true vs. predicted values with alignment to the diagonal indicating accuracy. The center plot displays residuals, highlighting error distribution centered around zero. The right plot visualizes true and predicted values across samples, with shaded areas representing error magnitudes.
@@
Figure D2: Evaluation of SVR performance for Layer best_layer on the atomic number. The left plot shows true vs. predicted values with alignment to the diagonal indicating accuracy. The center plot displays residuals, highlighting error distribution centered around zero. The right plot visualizes true and predicted values across samples, with shaded areas representing error magnitudes.
@@
Figure D3: Evaluation of SVR performance for Layer best_layer on the group. The left plot shows true vs. predicted values with alignment to the diagonal indicating accuracy. The center plot displays residuals, highlighting error distribution centered around zero. The right plot visualizes true and predicted values across samples, with shaded areas representing error magnitudes.
@@
Figure D4: Evaluation of SVR performance for Layer best_layer on the period. The left plot shows true vs. predicted values with alignment to the diagonal indicating accuracy. The center plot displays residuals, highlighting error distribution centered around zero. The right plot visualizes true and predicted values across samples, with shaded areas representing error magnitudes.
@@
Figure D5: Evaluation of SVR performance for Layer best_layer on the electronegativity. The left plot shows true vs. predicted values with alignment to the diagonal indicating accuracy. The center plot displays residuals, highlighting error distribution centered around zero. The right plot visualizes true and predicted values across samples, with shaded areas representing error magnitudes.

### v2

@@ E.2 Blessing of dimensionality
When the dimensionality is very high, the most of random vector pairs approach orthogonality. We illustrate this by sampling pairs of vectors in an 8129-dimensional space (corresponding to the residual stream vector dimension of Meta-Llama-3.1-70B) and computing their cosine similarities. The 99.9% confidence interval (CI) provides an estimate of the expected cosine similarity range at each dimensionality:
@@
Figure E.2: Cosine similarity distribution of random vector pairs in an 8129-dimensional space, with a 99.9% confidence interval ($-0.0364$, $0.0364$) shown by the dashed lines.
The specific distribution is shown in Fig. E.1, where the confidence interval is extremely narrow (±$\pm 0.036$), indicating that random vector pairs exhibit highly consistent cosine similarities. This suggests that the learned weights of the linear probe across different feature pairs are effectively uncorrelated, exhibiting only random alignment. In Fig. 4, the shaded region represents the 99.9% confidence interval for the cosine similarities of high-dimensional random vectors, further supporting this observation.
Appendix F Linear probing detailed results
In Sec.4.1, we applied linear probing to train linear-kernel Support Vector Regression (SVR) models for each layer $l$ and attribute $A_{j}$, using the residual stream dataset $H_{j}(l)$. The ground truth values $y_{i,j}$ correspond to the attribute values of each element $X_{i}$.
F.1 Why $R^{2}$ cannot reach 1
Even with a perfectly trained model and a sufficiently large dataset, achieving an $R^{2}$ of 1 in linear probing is impossible. The model’s output token ’1’ indicates that the residual stream at the last token position in the final layer leads to the highest logit for token 1’s ID after the final linear transformation. However, do token residual stream in the final layer of numbers exhibit a perfect linear relationship with their real numerical values? Token embeddings are learned representations that capture semantic relationships between tokens, but they are not guaranteed to align linearly with numerical values.
@@ F.1 Why $R^{2}$ cannot reach 1
F.2 Confusion matrix
Figure F.1: Confusion Matrix on categorical linear probing of attribute ‘Group’ on the middle layer.
F.3 Question prompts
We applied the Mann–Kendall test to $\Delta R^{2}$ in the upper 50% of layers for each model–attribute pair (15 in total). Benjamini–Hochberg adjustment (FDR = 0.05) retained significance in 12 pairs, with τ ranging from 0.36 to 0.94 (median 0.79; Table F.3).
Figure F.2: R² scores for linear probes trained on target properties and evaluated on representations from question and continuation prompts.
Table F.1: Mann–Kendall trend test on $\Delta R^{2}$ (difference between continuation and question prompt R²) within depth layer depths 0.5–1.0. $\tau$: Kendall’s tau (trend strength). (*) indicates significance after Benjamini–Hochberg FDR correction ($\alpha=0.05$).
Model Attribute $\tau$ $p$
Meta–Llama–3.1–70B Group 0.95 $<0.001$*
Period 0.84 $<0.001$*
Electronegativity 0.73 $<0.001$*
Atomic Mass 0.53 $<0.001$*
Atomic Number 0.41 $<0.001$*
Llama–2–7b–hf Electronegativity 0.80 $<0.001$*
Atomic Mass 0.80 $<0.001$*
Group 0.75 $<0.001$*
Atomic Number 0.55 $0.003$*
Period 0.40 $0.034$*
Llama–3.1–8B Group 0.57 $0.003$*
Electronegativity 0.52 $0.006$*
Atomic Number 0.20 $0.300$
Atomic Mass 0.20 $0.300$
Period -0.05 $0.822$
F.4 ‘Non-matching’ and ‘no mention’ prompts
F.4.1 Non-matching attribute pairs
To contrast explicit and implicit attribute cues, we selected a set of non-matching attribute pairs with minimal direct statistical dependency. Each pair was chosen to lack strong linear or monotonic relationships, and we verified on our 50-element subset that all selected pairs satisfy $|{\rm Pearson}\,r|<0.30$, $|{\rm Spearman}\,\rho|<0.30$, and $R^{2}<0.15$ (Table F.2). These thresholds ensure that no pair can be trivially recovered through simple linear or monotonic mappings.
Table F.2: Linear and monotonic correlations for non-matching attribute pairs (50-element subset).
Pair |Pearson r| |Spearman ρ| Linear R²
Group–Atomic Number 0.044 0.070 0.002
Group–Period 0.255 0.300 0.065
Group–Mass 0.037 0.066 0.001
Electronegativity–Atomic Number 0.154 0.038 0.024
Electronegativity–Mass 0.147 0.039 0.022
F.4.2 Detailed results
Figure F.3: R² score trends for ‘no mention’ cases. Regression linear probing on the element token residual stream with 5-fold cross-validation was performed on residual streams, and R2 scores on the test set are shown for each attribute.
Figure F.4: R² and $\Delta R^{2}$ scores for linear probes trained on target properties and evaluated on representations from matching and non-matching prompts. $\Delta R²$ is defined as R²(matching prompt) minus R²(non-matching prompt).
Figure F.5: Average $\Delta R^{2}$ across five attributes, with with 95% confidence interval shaded. $\Delta R^{2}=R^{2}_{\text{match}}-R^{2}_{\text{non-match}}$.
Table F.3: Mann–Kendall trend test on $\Delta R^{2}$ (difference between matching and non-matching prompt R²) within layer depths 0.6–1.0. $\tau$: Kendall’s tau (trend strength). “*” indicates significance after Benjamini–Hochberg FDR correction ($\alpha=0.05$).
Model Attribute Pair $\tau$ $p$
Meta–Llama–3.1–70B Electronegativity – Atomic Number $0.923$ $<0.001$*
Group – Atomic Number $0.911$ $<0.001$*
Electronegativity – Atomic Mass $0.895$ $<0.001$*
Group – Period $0.625$ $<0.001$*
Group – Atomic Mass $0.452$ $<0.001$*
Llama–2–7b–hf Group – Atomic Number $0.872$ $<0.001$*
Group – Atomic Mass $0.846$ $<0.001$*
Electronegativity – Atomic Mass $0.795$ $<0.001$*
Group – Period $0.769$ $<0.001$*
Electronegativity – Atomic Number $0.769$ $<0.001$*
Llama–3.1–8B Electronegativity – Atomic Mass $0.821$ $<0.001$*
Group – Period $0.769$ $<0.001$*
Group – Atomic Mass $0.744$ $<0.001$*
Electronegativity – Atomic Number $0.641$ $0.003$*
Group – Atomic Number $-0.154$ $0.502$
F.5 Detailed results of the best layer
@@
In the main text Sec. 4.1, we used SVR for linear regression probing. Figures F.6, F.7, F.8, F.9, and F.10 present the detailed $R^{2}$ performance of the best layer for each attribute—atomic number, atomic mass, electronegativity, period, and group—across three different models.
@@
Figure F.6: Evaluation of SVR performance for Layer best_layer on the atomic mass. The left plot shows true vs. predicted values with alignment to the diagonal indicating accuracy. The center plot displays residuals, highlighting error distribution centered around zero. The right plot visualizes true and predicted values across samples, with shaded areas representing error magnitudes.
@@
Figure F.7: Evaluation of SVR performance for Layer best_layer on the atomic number. The left plot shows true vs. predicted values with alignment to the diagonal indicating accuracy. The center plot displays residuals, highlighting error distribution centered around zero. The right plot visualizes true and predicted values across samples, with shaded areas representing error magnitudes.
@@
Figure F.8: Evaluation of SVR performance for Layer best_layer on the group. The left plot shows true vs. predicted values with alignment to the diagonal indicating accuracy. The center plot displays residuals, highlighting error distribution centered around zero. The right plot visualizes true and predicted values across samples, with shaded areas representing error magnitudes.
@@
Figure F.9: Evaluation of SVR performance for Layer best_layer on the period. The left plot shows true vs. predicted values with alignment to the diagonal indicating accuracy. The center plot displays residuals, highlighting error distribution centered around zero. The right plot visualizes true and predicted values across samples, with shaded areas representing error magnitudes.
@@
Figure F.10: Evaluation of SVR performance for Layer best_layer on the electronegativity. The left plot shows true vs. predicted values with alignment to the diagonal indicating accuracy. The center plot displays residuals, highlighting error distribution centered around zero. The right plot visualizes true and predicted values across samples, with shaded areas representing error magnitudes.